• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name E. Schulte

    Log entry time 15:55:57 on November 07, 2008

    Entry number 247369

    keyword=Day Shift Summary, November 7, 2008

    08:00 Shift starts with target motion problems. The target crew is
    attempting to reset the target mover.

    08:26 After target IOC reboot, the target moved to the "home" position
    under it's own power. Fortunately the safety system seemed to work.
    After several hair-raising moments, we went back to trying to get the
    beam positions centered up for the 3He target.

    08:29 Requested vertical move on beam to BPMA (1,1) BPMB (-1,1) for the
    3He cell.

    08:34 Reset positions are set and MCC is going to run an "F-Op" and try
    to get the FFB working.

    08:58 Move beam back to BPMA(1,4) BPMB(-1,4) for a half-hour carbon foil run.

    09:09 Request to move to optics target.

    09:10 Turned off BB wire chambers. We will request 5uA when we get that
    far. We may go to 10 uA depending upon the rates.

    09:14 Decided to request the Controlled Access now, since MCC is having
    problems with a magnet trim. We will go back to Kalyan's test when we
    come out of the access.

    09:26 The 3He target is spin up.

    09:39 Noticed that the target oven temperature has started to fall much
    below the set-point temperature of 230 (it's about 210 now).

    10:08 Target crew performs target motion test.

    10:39 Requested raster back on 4.5x4 (MCC units).

    12:26 Vincent & I worked on the BPM Raster problem. While we were gone,
    apparently HRS-L Q1 tripped. But it's resetting.

    13:34 During the shift checklist we found that BigBite was set to 200 A??
    Jack had to replace a UPS.

    13:39 Brad reset the BigBite magnet power supply.

    13:51 Now we're going to try the BPM Pedestal again.

    14:00 MCC to send tune beam S. Nanda will help MCC tune the beam through
    the Compton! Yea!

    14:18 MCC to send tune beam. We are currently in empty target.

    14:38 Changed to BeO target at MCC's request.

    15:20 MCC is going to send tune beam to fix up some orbit problems.

    15:55 MCC has been intermittently sending CW. They have not yet given
    any update as to what they are doing.